The firewall in the DTC is experiencing issues. This has led to networking issues in the DTC. Since many of our services depend on the network users may experience different issues that are associated with this. These include but are not limited to: Not being to access the internet, not being able to login to your machine, your machine "freezing", and, not being able to print. There may be other issues you are experiencing that are not listed explicitly here as well. We will keep you updated and are working on getting this issue fixed ASAP.
At around 8:20 this morning Keller Hall suffered a power outage. The power came back at around 9:20 but multiple systems are still affected. You may experience networking issues and/or login issues. We are working on getting all services back up and running ASAP. Thank you for your patience.
During the CSE-IT maintenance window Sunday morning, 5/20/2018, a firmaware update resulted in some CSE-IT Virtual Server infrastructure becoming unavailable. Services affected include CSE-IT supported websites which depend on a CSE-IT managed database, and Computer Science UNIX accounts.
Resolved 2:27 p.m. 5/20/2018
CSE-IT Major Incident In Resolved 04/23/18: At 12:22 pm, CSE-IT began receiving reports of issues regarding the DTC web service. Symptoms include: dtc.umn.edu being offline. We resolved the problem around 7:30 pm on Tuesday, 04/24/18.
Most OIT and CSE-IT systems were down for maintenance on Saturday and Sunday. CSE-IT Services that were affected included:
- CS e-mail
- CSE-IT Home Directories
- CSE-IT Project Space
- CSE-IT Databases and Websites
Maintenance has been completed.
Services that are stilll unavailable:
- CS Samba Drive
Please report any additional issues that persist to firstname.lastname@example.org
There was an issue with hosts in the DTC receiving DHCP addresses after the maintenance on Wednesday. The issue has been resolved and these hosts should be able to connect once more. Please contact email@example.com if connection issues persist.
Our mysql server has stopped working a couple of times today. We are actively working to determine the cause and come up with a solution. If you notice any issues please report them right away. At 10pm we will briefly shutdown the servers to apply a patch.
We are currently investigating issues with the Digital Technology Center network.